OpenStudy (geneticrockhopper247):

Please help. I am totally lost!!!!! Identify the vertex, focus, and directrix of the parabola with the equation x^2 – 6x – 8y + 49 = 0.

OpenStudy (anonymous):

put it into the form for a parabola.

OpenStudy (geneticrockhopper247):

In standard form, it would be (x-3)^2=8(y-5), right???

OpenStudy (geneticrockhopper247):

Putting it in vertex from is where I have trouble

OpenStudy (anonymous):

use this!

OpenStudy (geneticrockhopper247):

can you explain what p is???

OpenStudy (geneticrockhopper247):

I really would like to put it into vertex form, though (it makes more sense)

OpenStudy (anonymous):

did you look at the attachment? it explains everything.

OpenStudy (geneticrockhopper247):

I did, but it is a little confusing (to me). Is p in this one like 1/(4c) in vertex form??

OpenStudy (geneticrockhopper247):

@Spacelimbus can you help me???

OpenStudy (anonymous):

do you know how to "complete the square"??

OpenStudy (geneticrockhopper247):

@kausaralley I tried and got (x-3)^2=8(y-5) in standard form. What I really want to know is if y=1/8(x-3)^2-5 is correct vertex form.

OpenStudy (geneticrockhopper247):

@kausarsalley

OpenStudy (anonymous):

i don't think so you are supposed to get a value i am going to help you through the procedure if you want

OpenStudy (geneticrockhopper247):

okay

OpenStudy (geneticrockhopper247):

basically, I did this: x^2-6x+9-8y+49=9 x^2-6x+9=8y-49+9 (x-3)^2=8y-40 and then factored the right side

OpenStudy (geneticrockhopper247):

Please help me @phi @Directrix @experimentX

OpenStudy (geneticrockhopper247):

All I really want to know is how to put it into vertex form. I think I can do the rest from there.

OpenStudy (experimentx):

sorry ... was busy around.

OpenStudy (experimentx):

(x-3)^2=8(y-5)

OpenStudy (geneticrockhopper247):

how do I put it into vertex form????

OpenStudy (experimentx):

vertex ... 3, 5

OpenStudy (geneticrockhopper247):

Okay, but I want to make sure I understand this. Basically, how did you get that answer???

OpenStudy (geneticrockhopper247):

Sorry to nag.

OpenStudy (experimentx):

(x-3)^2=8(y-5) 3 5

OpenStudy (geneticrockhopper247):

The reason I wanted to use the vertex form is because I can find the directix and focus from it. How would you go about that with this??

OpenStudy (experimentx):

I don't know I almost forgot everything about conics .. let's see what I can do.

OpenStudy (experimentx):

 how do you find it's directrix?